OpenStudy (anonymous):

The sum of two consecutive odd integers is 28. What are the integers? Which let statement(s) would you use? (a) Let x = the integers (b) Let x = 1st integer. Let x + 2 = 2nd integer. (c) Let x = 1st integer. Let x + 1 = 2nd integer.

OpenStudy (anonymous):

B

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

Let 2x+1 be the first and 2x+3 the second.

OpenStudy (anonymous):

well, first let us get clarified a concept important to solve this numerical. "The sum of two consecutive odd integers is 28." Now the difference between two odd numbers is 2. So this means that one odd number has to be greater than the other by 2. let the smaller number be 'X' therefore the bigger number will be 'x+2' thus x + (x+2) = 28 therefore, x + x + 2 = 28 now we subtract "2" from both the sides... therefore, 2x = 26 now dividing throughout by 2 we get x = 13 therefore, x + 2 = 13+2 = 15 thereofre, the numbers are 13 AND 15

OpenStudy (anonymous):

Yes (b) is the correct answer as difference of two consecutive odd integers is 2

OpenStudy (anonymous):

Just for your info, the difference of two consecutive even integers is also 2

OpenStudy (anonymous):

I know